# Simulating the shaping of point-symmetric structures in the jittering jets explosion mechanism

### v2

@@ 3.2 Mimicking observations
    Usually bubbles are inflated by jets. However, in some cases small bubbles might appear between jet axes. These are the two opposite observed bubbles along $Z=0$ in the middle panel of Figure 7.
@@
    The upper panel of Figure 8 shows a morphology of three pairs of bubbles. These are jet-inflated bubbles. The radioactive decay of nickel can also form bubbles (nickel bubbles, e.g., Orlando et al. (2025)). The decay of nickel and inflation of bubbles occurs long after the explosion. Nickel bubbles will form as two opposite bubbles if a pair of jets ejected two opposite clumps of nickel.
The six simulations show the rich variety of point-symmetric morphologies that jets can shape, some resembling observed CCSN remnants, although we were not aiming at any specific observation.
@@ 4 Results: Rayleigh-Taylor instabilities and vortices
As a jet deposits energy into the jet-core interaction region, it forms a high-pressure volume that expands. As a result of this expansion, the density in the bubble becomes lower than that of its surroundings, while its pressure stays higher. The zones of opposite pressure and density gradients are prone to Rayleigh-Taylor instability (RTI). To locate these zones and their growth rate, we refer to the quantity
@@
Where $\rho$ is the density, $P$ is the pressure and ${\rm sgn}(\vec{\nabla}P\cdot\vec{\nabla}\rho)$ is the sign of the product of pressure and density gradients. If $f_{st}$ is negative, the region is unstable with a typical growth rate of $-f_{st}$, and a typical growth time of $-1/f_{st}$.
The upper row panels of Figure 9 present maps of $f_{\rm st}$ in units of ${~{}\rm s}^{-1}$ for simulation A3 at $t=1.36{~{}\rm s}$ in three planes of the grid. In the deep red zones $f_{\rm st}>0$ and the red areas are RTI stable. The yellow to deep blue zones are RTI unstable, with the fastest growing rate in the deep blue zones, a growth time of about 1 second. Although we present the results at the last time of the simulation, the RTI exists along the entire jet-core interaction. The growth time is shorter than the simulation time, and the RTI has time to grow to the non-linear regime. The RTI explains the breakup of clumps to knots (smaller clumps; Section 3). The unstable zones are thin surfaces that appear as thin filaments on the planes. The unstable surfaces occupy a small fraction of the volume. As with water laying on oil, the unstable zones are near the interface between the two media and might have a small filling factor, but they still cause the development of large-amplitude instabilities. Therefore, even if the RTI unstable zones with short time growth have a small filling factor, the RTI influences the small-scale structure of the ejecta, e.g., breaking clumps and filaments, because the RTI growth time is shorter than the simualtion time. The panels in the upper row of Figure 10 present the RTI maps of simulation A5 at $t=1.94{~{}\rm s}$.
Figure 9: RTI growth rate and vorticity color maps of simulation A3 at $t=1.36{~{}\rm s}$ in the planes, from left to right, $z=0$, $x=0$, and $y=0$. The upper panels show $f_{st}$, calculated using equation (2). Deep-red areas are RTI stable, while others are unstable with a typical growth time of $-1/f_{st}$. Color bar is from $-1{~{}\rm s}^{-1}$ (deep blue) to $\geq 0$ (deep red). Lower panels show the vorticity component perpendicular to the given plane, (×$(\nabla\times\vec{v})_{\perp}$. The color bar runs from $-1{~{}\rm s}^{-1}$ (deep blue clockwise) to $1{~{}\rm s}^{-1}$ (deep red; counterclockwise).
@@
In the RTI, low-density fingers protrude into the high-density low-pressure zone and high-density fingers protrude into the low-density high-pressure zone. The result is the formation of vortices on small scales in RTI unstable zones. On the other hand, the interaction of the jets with the core material forms mainly vortices on large scales, i.e., of the order of the scales of the bubbles the jets inflate. This difference is best seen in the middle panels of Figure 9. The RTI unstable zones that are zones with blue stripes on the upper-middle panel of Figure 9, have red-blue stripes in the middle-lower panel that show the vorticity component perpendicular to the plane, i.e., (×$(\nabla\times\vec{v})_{\perp}$. These are small-scale vortices. The continuous large blue and red areas in the middle-lower panel are large-scale vortices induced by the backflow that the jets form; comparing the two middle panels, shows these areas to be RTI stable. The role of the instabilities and vortices in nucleosynthesis during the explosion process and the mixing of elements are the topics of a future study in this series of studies. The lower rows of Figures 9 and 10 show the vortex values in the other planes of simulation A3 and A5.
In Figure 11 we present velocity maps in two planes of simulation A3 and A5 to further demonstrate the large-scale vortices that the jets induce. The colors depict the magnitude of the velocity (red being fast), and the arrows are along the flow direction. In the lobes, the highest outward velocity is along the symmetry axis. The expansion velocity is lower on the sides of the symmetry axis of a lobe/bubble. In the frame of the expanding lobes, there are large vortices with a flow from the center out and then back, i.e., a backflow.
@@
The velocity maps of Figure 11 show that near the center the material flows inward. We do not treat the center accurately in this study. Firstly, the newly born NS has a wind that can push material out. Secondly, and more importantly, the backflow near the center might lead to further accretion that launches more jets. In a future study, we will include later jets and the expansion of the forward shock out from the core.
The velocity maps of simulation A3 in the upper panels of Figure 11 show pairs of expanding lobes. The shocked core’s material occupies the outer volume of these lobes, while the original jets’ material is inside these lobes. To demonstrate this, we present in Figure 12 the tracer of the jets of simulation A3 in the same planes as the velocity maps. The tracer is a numerical quantity that traces the flow of a designated volume. In the present case, it is the material of the jets. At jets’ launching the tracer in the jet equals 1, i.e., pure jets’ material. With time, the value of the tracer in a cell shows the fraction of mass that started in the jets; the tracer can have any value from 0 to 1. The size of the lobes in the tracer maps of Figure 12 is much smaller than the expanding high-velocity lobes of the upper panels of Figure 11. This shows that the original jets’ material is deep inside the fast-expanding lobes.
@@ 5 Summary
To explain the large fraction of point-symmetric CCSN remnants, we conducted six 3D hydrodynamical simulations of the JJEM. In these simulations, we launch several pairs of jittering jets in varying directions inside a collapsing massive stellar core model. We examined the morphologies that the jittering jets shape. In this study, we do not follow the backflow or the NS wind, so we did not analyze the grid’s inner $\simeq 15\%$.
@@
We summarized the results of our simulations in a list of points in Section 3.2. We consider most significant our finding that on average a pair of jets might form more than one pair of clumps, that small bubbles might form not only along the jets’ axis, that RTI might break clumps into smaller clumps (knots), and that jets can form an ‘H-shaped’ CCSN remnant like W49B. These are significant results because they explain specific observed properties of some CCSN remnants. Our study further solidifies the JJEM as the primary explosion mechanism of CCSNe.
@@
In future studies we will follow the explosion to break out from the star, the launching of late jets due to the backflow (see Figure 11 for the flow structure), the nucleosynthesis as the jets interact with the core material, and the role of vortices and RTIs (Figures 9 and 10) in distributing and mixing the newly synthesis elements. Our results will lead the search for more jet signatures in CCSNe and CCSN remnants.
@@
Words are in place here on the general approach of our study, which is very common in studying active galactic nuclei (AGN) feedback and in shaping planetary nebulae. It is based on these ingredients. (1) Strong observational indications of jet activity at present or in the past, like presently active jets, outcome of past jets (like pairs of bubbles and lobes), or gamma rays in gamma ray bursts. We have the point symmetric morphologies of more than ten CCSN remnants (Section 1), and more CCSN remnants with one pair of ears, that indicate jet activity. (2) The presence of an accreting compact object, like a super-massive black hole in an AGN, a companion to the progenitor of a planetary nebula, or a stellar mass black hole in gamma ray bursts. CCSNe form an NS or a black hole at the center that accretes more core material. (3) Manual jet launching without resolving the compact object’s vicinity. Namely, the jets are launched from a distance much larger than the accretion disk that launches them. The numerical launching of jets does not refer to the physical mechanism by which the accretion disk launches the jets, which is a difficult and unsolved process. Particularly relevant to us, a recent study finds that present numerical simulations do not have the resolution to correctly follow magnetic field reconnection events that are crucial for jet launching in CCSNe (Soker, 2024b).
@@
Identifying point-symmetric morphologies in the images of CCSN remnants in the last two years (Section 1) has been a breakthrough in our understanding of the explosion mechanism of CCSNe. By reproducing the general point-symmetric structures observed in these CCSN remnants, our study adds to and solidifies the growing observational evidence that the JJEM is the primary explosion mechanism of CCSNe.
@@ Acknowledgements
We thank Dmitry Shishkin for helpful comments and discussions. Amir Michaelis is supported by the European Union’s Horizon 2020 research and innovation program under grant agreement No 865932-ERC-SNeX. We acknowledge the Ariel HPC Center at Ariel University for providing computing resources that have contributed to the research results reported in this paper.

### v3

@@ 3.2 Mimicking observations
    Usually, bubbles are inflated by jets. However, in some cases, small bubbles might appear between jet axes. These are the two opposite observed bubbles along $Z=0$ in the middle panel of Figure 7.
@@
    The upper panel of Figure 8 shows a morphology of three pairs of bubbles. These are jet-inflated bubbles. The radioactive decay of nickel can also form bubbles (nickel bubbles, e.g., Orlando et al. (2025)). The decay of nickel and the inflation of bubbles occur long after the explosion. Nickel bubbles will form as two opposite bubbles if a pair of jets ejects two opposite clumps of nickel.
The six simulations demonstrate the rich variety of point-symmetric morphologies that jets can shape, some of which resemble observed CCSN remnants, although we were not aiming to replicate any specific observation.
@@ 4 Results: Rayleigh-Taylor instabilities and vortices
As a jet deposits energy into the jet-core interaction region, it forms a high-pressure volume that expands. As a result of this expansion, the density within the bubble decreases below that of its surroundings, while its pressure remains higher. Zones of opposite pressure and density gradients are prone to the Rayleigh-Taylor instability (RTI). To locate these zones and their growth rate, we refer to the quantity
@@
Where $\rho$ is the density, $P$ is the pressure, and ${\rm sgn}(\vec{\nabla}P\cdot\vec{\nabla}\rho)$ is the sign of the product of pressure and density gradients. If $f_{st}$ is negative, the region is unstable with a typical growth rate of $-f_{st}$, and a typical growth time of $-1/f_{st}$.
The upper row panels of Figure 9 present maps of $f_{\rm st}$ in units of ${~{}\rm s}^{-1}$ for simulation A3 at $t=1.36{~{}\rm s}$ in three planes of the grid. In the deep red zones $f_{\rm st}>0$ and the red areas are RTI stable. The yellow to deep blue zones are RTI unstable, with the fastest growing rate in the deep blue zones, a growth time of about 1 second. Although we present the results at the final time of the simulation, the RTI persists throughout the entire jet-core interaction. The growth time is shorter than the simulation time, allowing the RTI to transition into the non-linear regime. The RTI explains the breakup of clumps to knots (smaller clumps; Section 3). The unstable zones are thin surfaces that appear as thin filaments on the planes. The unstable surfaces occupy a small fraction of the volume. As with water lying on oil, unstable zones are located near the interface between the two media and may have a small filling factor, but they still contribute to the development of large-amplitude instabilities. Therefore, even if the RTI unstable zones with short time growth have a small filling factor, the RTI influences the small-scale structure of the ejecta, e.g., breaking clumps and filaments, because the RTI growth time is shorter than the simulation time. The panels in the upper row of Figure 10 present the RTI maps of simulation A5 at $t=1.94{~{}\rm s}$.
Figure 9: RTI growth rate and vorticity color maps of simulation A3 at $t=1.36{~{}\rm s}$ in the planes, from left to right, $z=0$, $x=0$, and $y=0$. The upper panels show $f_{st}$, calculated using equation (2). Deep-red areas are RTI stable, while others are unstable with a typical growth time of $-1/f_{st}$. Color bar is from $-1{~{}\rm s}^{-1}$ (deep blue) to $\geq 0$ (deep red). The lower panels show the vorticity component perpendicular to the given plane, (×$(\nabla\times\vec{v})_{\perp}$. The color bar runs from $-1{~{}\rm s}^{-1}$ (deep blue clockwise) to $1{~{}\rm s}^{-1}$ (deep red; counterclockwise).
@@
In the RTI, low-density fingers protrude into the high-density low-pressure zone, and high-density fingers protrude into the low-density high-pressure zone. The result is the formation of vortices on small scales in RTI unstable zones. On the other hand, the interaction of the jets with the core material forms mainly vortices on large scales, i.e., of the order of the scales of the bubbles the jets inflate. This difference is best seen in the middle panels of Figure 9. The RTI unstable zones that are zones with blue stripes on the upper-middle panel of Figure 9, have red-blue stripes in the middle-lower panel that show the vorticity component perpendicular to the plane, i.e., (×$(\nabla\times\vec{v})_{\perp}$. These are small-scale vortices. The continuous large blue and red areas in the middle-lower panel are large-scale vortices induced by the backflow that the jets form; comparing the two middle panels, shows these areas to be RTI stable. The role of instabilities and vortices in nucleosynthesis during the explosion process and the mixing of elements are topics for a future study in this series. The lower rows of Figures 9 and 10 show the vortex values in the other planes of simulation A3 and A5.
In Figure 11 we present velocity maps in two planes of simulation A3 and A5 to further demonstrate the large-scale vortices that the jets induce. The colors represent the magnitude of the velocity (red indicating high velocity), and the arrows are aligned with the flow direction. In the lobes, the highest outward velocity is along the axis of symmetry. The expansion velocity is lower on the sides of the symmetry axis of a lobe/bubble. In the frame of the expanding lobes, there are large vortices with a flow from the center out and then back, i.e., a backflow.
@@
The velocity maps of Figure 11 show that near the center, the material flows inward. We do not treat the center accurately in this study. Firstly, the newly born NS has a wind that can push material out. Secondly, and more importantly, the backflow near the center may lead to further accretion, which in turn launches more jets. In a future study, we will include later jets and the expansion of the forward shock out from the core.
The velocity maps of simulation A3 in the upper panels of Figure 11 show pairs of expanding lobes. The shocked core’s material occupies the outer volume of these lobes, while the original jets’ material is inside these lobes. To demonstrate this, we present in Figure 12 the tracer of the jets of simulation A3 in the same planes as the velocity maps. The tracer is a numerical quantity that traces the flow of a designated volume. In the present case, it is the material of the jets. At jets’ launching, the tracer in the jet equals 1, i.e., pure jets’ material. Over time, the value of the tracer in a cell indicates the fraction of mass that originated from the jets; the tracer can assume any value between 0 and 1. The size of the lobes in the tracer maps of Figure 12 is much smaller than the expanding high-velocity lobes of the upper panels of Figure 11. This indicates that the original jets’ material is located deep within the rapidly expanding lobes.
@@ 5 Summary
To explain about a dozen point-symmetric CCSN remnants, we conducted six 3D hydrodynamical simulations of the JJEM. In these simulations, we launch several pairs of jittering jets in varying directions inside a collapsing massive stellar core model. We examined the morphologies that the jittering jets shape. In this study, we do not follow the backflow or the NS wind, so we did not analyze the grid’s inner $\simeq 15\%$.
@@
We summarized the results of our simulations in a list of points in Section 3.2. We consider most significant our finding that on average a pair of jets might form more than one pair of clumps, that small bubbles might form not only along the jets’ axis, that RTI might break clumps into smaller clumps (knots), and that jets can form an ‘H-shaped’ CCSN remnant like W49B. These are significant results because they explain specific observed properties of some CCSN remnants. Our study further supports the JJEM as a viable mechanism of CCSNe.
@@
In future studies we will follow the explosion to break out from the star, the launching of late jets due to the backflow (see Figure 11 for the flow structure), the nucleosynthesis as the jets interact with the core material, and the role of vortices and RTIs (Figures 9 and 10) in distributing and mixing the newly synthesis elements. Our results will lead the search for more jet signatures in CCSNe and CCSN remnants. For example, we expect that when a jet breaks out from the star, possibly late jets, it will experience a shock and, for the lack of dense ambient gas, will lose collimation. This is observed in jets in galaxies and planetary nebulae. In CCSNRs, the outer uncollimated jet’s material is referred to as a blowout, and was studied in the frame of the JJEM for the Cygnus Loop and SNR G309.2–00.6 by Shishkin et al. (2024), and for SNR G0.9+0.1 by Soker (2025).
@@
Words are in place here on the general approach of our study, which is very common in studying active galactic nuclei (AGN) feedback and in shaping planetary nebulae. It is based on these ingredients. (1) Strong observational indications of jet activity at present or in the past, like presently active jets, outcome of past jets (like pairs of bubbles and lobes), or gamma rays in gamma ray bursts. We have the point-symmetric morphologies of more than ten CCSN remnants (Section 1), and more CCSN remnants with one pair of ears, which indicate jet activity. (2) The presence of an accreting compact object, like a super-massive black hole in an AGN, a companion to the progenitor of a planetary nebula, or a stellar mass black hole in gamma ray bursts. CCSNe form an NS or a black hole at the center that accretes more core material. (3) Manual jet launching without resolving the compact object’s vicinity. Namely, the jets are launched from a distance much larger than the accretion disk that launches them. The numerical launching of jets does not refer to the physical mechanism by which the accretion disk launches the jets, which is a difficult and unsolved process. Particularly relevant to us, a recent study finds that current numerical simulations lack the resolution to accurately capture magnetic field reconnection events, which are crucial for jet launching in CCSNe (Soker, 2024b).
@@
Identifying point-symmetric morphologies in the images of CCSN remnants in the last two years (Section 1) has been a breakthrough in our understanding of the explosion mechanism of CCSNe. By reproducing the general point-symmetric structures observed in these CCSN remnants, our study encourages further simulations of the JJEM aiming at reproducing more morphologies and other CCSN properties, e.g., distributions of radioactive isotopes.
@@ Acknowledgements
We thank Dmitry Shishkin for helpful comments and discussions, and an anonymous referee for helpful comments. Amir Michaelis is supported by the European Union’s Horizon 2020 research and innovation program under grant agreement No 865932-ERC-SNeX. We acknowledge the Ariel HPC Center at Ariel University for providing computing resources that have contributed to the research results reported in this paper.